Fiona is seven years older than Gavin. Alex is 12 years younger than Gavin. Is the total of the ages of 73 how old is the youngest of them?
Answer by ramkikk66(644)

Let Gavin's age be x
Then Fiona's age = x + 7
Then Alex's age = x - 12
Sum of the ages = x + x + 7 + x - 12 = 73 (given)
i.e. 3*x - 5 = 73
i.e. 3*x = 73 + 5 = 78
Solving for x, x = 78 / 3 = 26
So Gavin's age = 26
Fiona's age = 26 + 7 = 33
Alex's age = 26 - 12 = 14
Confirming, 26 +33 + 14 = 73
Youngest person is Alex whose age is 14